OpenStudy (anonymous):

Determine the equation of the line, in standard form, that will get your spacecraft from the (1,2) to95 ,5)

OpenStudy (anonymous):

* (1,2) to (5,5)

OpenStudy (anonymous):

I am stuck. So far i have the slope 3/4 and i was using the point (5, 5). so far i have y-5= 3/4(x -5), i feel like i always make a mistake after this. can someone help?

OpenStudy (anonymous):

@thomaster @myininaya @hartnn

OpenStudy (anonymous):

i'm sorry to bother you guys, i'm new here...

hartnn (hartnn):

y-5= 3/4(x -5) is absolutely correct! with correct method :)

hartnn (hartnn):

so in which form do u need the equation?

hartnn (hartnn):

i mean u know what standard form is?

OpenStudy (anonymous):

i need standard form which is Ax + By = C

hartnn (hartnn):

right, so for that lets multiply 4 on both sides first, what u get?

OpenStudy (anonymous):

multiply 4? where is the 4? isn't it multiply 3/4 to (x-5)

OpenStudy (anonymous):

? i'm not sure

hartnn (hartnn):

it is, now we have a fraction there , the 3/4 how to eliminate that? just multiply 4 to each term of the equation so, 3/4*4=3 <----not a fraction makes sense?

OpenStudy (anonymous):

is that ok? well if we do that so far we would have y - 5 = 3 (x-5) which would be y-5= 3x -15.. then adding 5 to both sides it would be y= 3x -10 then moving 3x to the left we would get 3x + y = -10? right i feel that there should be a fraction in the front?

hartnn (hartnn):

i said , multiply 4 to each term :) so, 4y-4*5=3(x-5) got this?

OpenStudy (anonymous):

ahhhh! you see! i make this kind of mistake always! so 4y - 20 = 3x - 15 would work too?

hartnn (hartnn):

in the standard form, we need constants on one side! so lets first add 20 to both sides what u get?

OpenStudy (anonymous):

4y = 3x +5?

hartnn (hartnn):

correct! lastly, just bring 3x on left, how ?

OpenStudy (anonymous):

ahh yes! but is that ok? 3x + 4y = 5, i feel like there should be a fraction? but if this is ok then i can always do my work like this right? i always get confused with fractions..

OpenStudy (anonymous):

by the way, thank you very very much! you helped me realized that there is more then one way or solving these types of problems :)

hartnn (hartnn):

you just can not move it to left side! you need to subtract 3x from both sides, so that on right side, 3x-3x =0 and 3x vanishs from right and -3x appers on left, makes sense ?

hartnn (hartnn):

thank me after you get the correct answer :)

OpenStudy (anonymous):

ahhh again, i did it again!! ahh -3x +4y =5 so i would have to multiply everything by -1 right?

OpenStudy (anonymous):

so if it is like that it would be (-1)-3x + (-1)4y= (-1)5, which is 3x -4y = -5?

hartnn (hartnn):

you can if you want to, but -3x +4y =5 is also acceptable

hartnn (hartnn):

but the convention is to write the co-efficient of x as positive, so yes, go ahead and do that

hartnn (hartnn):

oh, u already did :P yes, 3x-4y=-5 is correct :D

hartnn (hartnn):

any doubts ?

hartnn (hartnn):

and since u r new here, \[ \begin{array}l\color{red}{\text{W}}\color{orange}{\text{E}}\color{#e6e600}{\text{L}}\color{green}{\text{C}}\color{blue}{\text{O}}\color{purple}{\text{M}}\color{purple}{\text{E}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#e6e600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{O}}\color{purple}{\text{p}}\color{purple}{\text{e}}\color{red}{\text{n}}\color{orange}{\text{S}}\color{#e6e600}{\text{t}}\color{green}{\text{u}}\color{blue}{\text{d}}\color{purple}{\text{y}}\color{purple}{\text{!}}\color{red}{\text{!}}\color{orange}{\text{}}\end{array} \]

OpenStudy (anonymous):

no doubts :) thank you very much... again haha :D

OpenStudy (anonymous):

and thank you for the welcome :D

hartnn (hartnn):

welcome ^_^

hartnn (hartnn):

i taught you math, will u teach me tennis ?

hartnn (hartnn):

gone? :(